# TSCLIP: Robust CLIP Fine-Tuning for Worldwide Cross-Regional Traffic Sign Recognition

### v1

@@ Abstract
Traffic sign is a critical map feature for navigation and traffic control. Nevertheless, current methods for traffic sign recognition rely on traditional deep learning models, which typically suffer from significant performance degradation considering the variations in data distribution across different regions. In this paper, we propose TSCLIP, a robust fine-tuning approach with the contrastive language-image pre-training (CLIP) model for worldwide cross-regional traffic sign recognition. We first curate a cross-regional traffic sign benchmark dataset by combining data from ten different sources. Then, we propose a prompt engineering scheme tailored to the characteristics of traffic signs, which involves specific scene descriptions and corresponding rules to generate targeted text descriptions for optimizing the model training process. During the TSCLIP fine-tuning process, we implement adaptive dynamic weight ensembling (ADWE) to seamlessly incorporate outcomes from each training iteration with the zero-shot CLIP model. This approach ensures that the model retains its ability to generalize while acquiring new knowledge about traffic signs. Our method surpasses conventional classification benchmark models in cross-regional traffic sign evaluations, and it achieves state-of-the-art performance compared to existing CLIP fine-tuning techniques (Fig. 1). To the best knowledge of authors, TSCLIP is the first contrastive language-image model used for the worldwide cross-regional traffic sign recognition task. The project website is available at: https://github.com/guoyangzhao/TSCLIP.
@@ I Introduction
Traffic sign recognition is a critical perceptual task in autonomous and assisted driving systems[1]. Traffic signs provide rich map features and road navigation information, which are crucial for driving safety and understanding the current scene[2]. Traditional traffic sign classification methods mainly rely on manually designed and extracted features such as color or shape, and use parameter-based classifiers for recognition [3]. These methods are heavily dependent on parameter tuning and feature design, making them susceptible to environmental factors and varying scenarios, resulting in lower recognition robustness.
In recent years, with the development of deep learning technology, convolutional neural networks (CNNs) have achieved automatic feature extraction and learning in high-dimensional spaces[4, 5], significantly reducing the difficulty of feature design and improving recognition performance. Using deep learning methods, high accuracy results have been achieved in the field of traffic sign recognition, far surpassing traditional feature design methods[6]. However, since CNN models are trained only on their respective datasets, their performance significantly deteriorates when tested on datasets from different environments or regions, even for the same categories[7, 8]. Furthermore, some datasets[9, 10, 6, 11] use only symbols to represent categories. While this has little impact on general classification tasks, it has a substantial impact on driving tasks that rely on the semantic information of different traffic signs to understand the environment.
@@
The aforementioned issues have gradually been addressed with the introduction of contrastive language-image pre-training (CLIP) [12]. CLIP, by pre-training on a large-scale image-text dataset, has demonstrated the ability to learn both general and distinctive visual representations through language supervision. By using independent encoders to extract features from input images and texts, CLIP aligns paired features in the same feature space and employs contrastive loss to formulate the learning objective[13]. CLIP has proven to exhibit excellent zero-shot performance in visual representation, enabling the recognition and classification of new data containing unseen category images in downstream tasks, showcasing strong recognition capabilities[14]. In the task of traffic sign recognition, leveraging CLIP’s powerful visual-language learning capability can enhance the understanding of sign semantic information, making it possible to achieve cross-region and environment recognition.
@@
In this work, to meet the requirements of cross-regional recognition, we first extract 46 mainstream and universal traffic sign categories from 10 existing datasets, creating a cross-regional traffic sign (CRTS) dataset. Based on the categories and feature distributions of traffic signs, we propose a prompt engineering scheme specifically designed for traffic signs. Regarding the TSCLIP model, we perform fine-tuning on the pre-trained CLIP model and introduce an adaptive dynamic weight ensembling (ADWE) fine-tuning scheme. Specifically, we dynamically integrate the results of each training iteration with the CLIP zero-shot model using adaptive factors. Ensuring the model maintains the generalization capabilities of the zero-shot model while learning new traffic sign knowledge.
Our primary contributions are as follows:
@@
    We develop TSCLIP, the first comparative language-image model designed for traffic sign recognition, achieving state-of-the-art cross-region testing performance and outperforming mainstream benchmark models.
@@ II-A Traffic Sign Recognition
The advantage of deep learning methods lies in their ability to automatically learn and extract complex features from data, leading to higher accuracy. [22] evaluated various CNNs and vision transformer models, showing that CNNs perform better in traffic sign classification. [23] proposed multi-scale CNN and multi-column CNN approaches that performed well in multiple datasets. [24] proposed a novel method using a limited common image set for traffic sign recognition, demonstrating excellent accuracy. [25] developed a lightweight CNN model that achieved near-perfect accuracy on the GTSRB dataset. [26] proposed a CNN model that achieved over 91% accuracy on an Indian dataset. Although deep learning methods can automatically learn features of different signs, their performance is often excellent only on the training dataset. Once transferred to unseen scenarios, the robustness and generalization significantly decline.
@@ II-B3 Full Fine-Tuning (FFT)
FFT updates all parameters of the pre-trained model to adapt to specific tasks. The optimization objective involves updating the weights of both the image and text encoders, with the loss function defined as:
@@ III-A1 Dataset Construction
In this study, we created a CRTS joint dataset based on mainstream open-source traffic sign datasets. To ensure regional diversity, we selected datasets from 10 different countries or regions[2, 3, 6, 9, 10, 11, 28, 27, 29, 30]. We extracted 46 commonly used categories by analyzing the distribution of similar categories and traffic sign regulations[31] across these datasets. Every dataset was cleaned, and all categories were standardized with corresponding semantic names.
@@ III-A2 Difference of Cross-Regional Traffic Signs
Fig. 2 shows examples of four traffic sign categories (No Overtaking, No Parking, No Pedestrians, and Stop) in different regional contexts. Traffic sign patterns vary across countries and regions, influenced by local culture and traffic regulations. Some regions have unique sign patterns, such as the No Overtaking signs in China and America. Additionally, some signs incorporate local languages, as seen in the Stop signs from China, Iran, India, Turkey, and Russia. Therefore, cross-regional traffic sign recognition poses a significant challenge, requiring models to handle continuously changing sign patterns.
@@ III-B2 Refinement of Scenario Descriptions ($\mathbf{S_{i}}$)
The scenario descriptions ($\mathbf{S_{i}}$) encompasses four critical elements (a-d):
@@ III-B3 Traffic Sign Category and Rules ($\mathbf{T_{i}}$)
By following this structured approach, we create $n$ diverse and dynamic prompt templates $\{\mathbf{P_{1}},\mathbf{P_{2}},\ldots,\mathbf{P_{n}}\}$, ensuring a comprehensive representation of traffic signs in various scenarios. This method enhances the model’s ability to understand and classify traffic signs by providing rich contextual information and explicit traffic rules, leading to more accurate and robust recognition.
@@ III-C TSCLIP Fine-Tuning Implementation
The core aim of the TSCLIP model’s robust fine-tuning framework (Fig. 3) is to combine the excellent generalization ability of the zero-shot model across all data distributions with the recognition capability of the fine-tuned model on traffic sign training data. This ensures the best of both worlds for specialized tasks and cross-regional generalization. Specifically, the fine-tuning framework consists of two main parts. Fig. 3(a) shows the initial fine-tuning of the TSCLIP model using image data and corresponding prompts from the training samples, ensuring the fine-tuned model learns domain-specific knowledge of traffic signs. Fig. 3(b) illustrates the adaptive dynamic weight ensembling of the latest training weights with the CLIP zero-shot model at the end of each training epoch, ensuring that the model retains a certain level of zero-shot generalization. Additionally, the robustness gains achieved by TSCLIP during the fine-tuning process do not incur extra computational costs during fine-tuning or inference.
@@ III-D1 Weight Ensembling
Interpolating model parameters is a classic idea in convex optimization[32]. Previous studies have shown that interpolation in the weight space can improve performance when models share part of the optimization trajectory [33]. Wise-FT is the first empirical study to explore the interpolation of non-convex models, specifically CLIP models, from the perspective of distributional robustness. Fig. 3(c) illustrates Wise-FT’s ensembling method, which involves weight ensembling with the zero-shot model after the entire fine-tuning of the CLIP model is completed. This weight ensembling method enhances robustness by forcibly injecting the weights of the zero-shot model into the final training results. Although experimental results show significant improvements, this approach overlooks the dynamic nature of CLIP model weights during the fine-tuning process.
@@ III-D3 Adaptive Factor
In our dynamic weight ensembling method, since the weights obtained from each epoch are dynamically changing and the ensembled weights are further iteratively trained, a fixed mixing coefficient is not sufficient to handle the dynamic nature of the weights. The mixing coefficient must adapt dynamically to the ever-changing training.
@@
This formula ensures that the ensemble proportion of the zero-shot model weights decreases smoothly over time while dynamically adjusting based on the relative performance of the trained model and the zero-shot model. This approach effectively balances the generalization and task-specific learning needs, leading to robust fine-tuning of the CLIP model for traffic sign recognition across different regions.
@@ IV-A Experiment Setup
All models in this study were trained in the PyTorch framework using the NVIDIA A100-PCIE-40GB GPUs. For training parameters, the CLIP-based models were set with a batch size of 512 and trained for 10 epochs with a learning rate of 0.00003. The classic classification model was trained based on the existing pre-trained model with a batch size of 128, trained for 100 epochs, and the learning rate was 0.0001.
@@ V Conclusion
To address the challenge of robust traffic sign recognition across regions and data distributions worldwide, we proposed the TSCLIP model and a CRTS benchmark dataset. Then, we developed a prompt engineering scheme that includes specific scene descriptions and corresponding rules, specifically made for traffic sign. The proposed ADWE method effectively combines fine-tuning model with zero-shot model, ensuring generalization to other environments while learning new traffic sign knowledge. In extensive cross-regional tests, TSCLIP significantly outperformed mainstream benchmark methods and achieved SOTA result compared to existing robust fine-tuning methods. The ablation experiments and visual analyses further validated and illustrated the effectiveness of our approach. Future research will involve collecting traffic sign image data worldwide to build the foundation model, enabling general recognition of traffic signs across regions.

### v2

@@ Abstract
Traffic sign is a critical map feature for navigation and traffic control. Nevertheless, current methods for traffic sign recognition rely on traditional deep learning models, which typically suffer from significant performance degradation considering the variations in data distribution across different regions. In this paper, we propose TSCLIP, a robust fine-tuning approach with the contrastive language-image pre-training (CLIP) model for worldwide cross-regional traffic sign recognition. We first curate a cross-regional traffic sign benchmark dataset by combining data from ten different sources. Then, we propose a prompt engineering scheme tailored to the characteristics of traffic signs, which involves specific scene descriptions and corresponding rules to generate targeted text descriptions. During the TSCLIP fine-tuning process, we implement adaptive dynamic weight ensembling (ADWE) to seamlessly incorporate outcomes from each training iteration with the zero-shot CLIP model. This approach ensures that the model retains its ability to generalize while acquiring new knowledge about traffic signs. To the best knowledge of authors, TSCLIP is the first contrastive language-image model used for the worldwide cross-regional traffic sign recognition task. The project website is available at: https://github.com/guoyangzhao/TSCLIP.
@@ I Introduction
Traffic sign recognition is a critical perceptual task in autonomous and assisted driving systems[1]. Traffic signs provide rich map features and road navigation information, which are crucial for driving safety and understanding the current scene[2]. Traditional traffic sign classification methods mainly rely on manually designed and extracted features such as color or shape, and use parameter-based classifiers for recognition [3]. These methods are heavily dependent on feature parameter tuning, making them susceptible to varying scenarios, resulting in lower recognition robustness.
In recent years, convolutional neural networks (CNNs) have achieved automatic feature extraction and learning in high-dimensional spaces[4, 5], significantly reducing the difficulty of feature design and improving recognition performance. Using deep learning methods, high accuracy results have been achieved in the field of traffic sign recognition, far surpassing traditional feature design methods[6]. However, since CNNs are trained only on their respective datasets, their performance significantly deteriorates when tested on datasets from different environments or regions, even for the same categories[7, 8]. Furthermore, some datasets[9, 10, 6, 11] use only symbols to represent categories. While this has little impact on general classification tasks, it has a substantial impact on driving tasks that rely on the semantic information of different signs to understand the environment.
@@
The aforementioned issues have gradually been addressed with the introduction of contrastive language-image pre-training (CLIP) [12]. By using independent encoders to extract features from input images and texts, CLIP aligns paired features in the same feature space and employs contrastive loss to formulate the learning objective[13]. CLIP has proven to exhibit excellent zero-shot performance in visual representation, enabling the recognition and classification of new data containing unseen category images in downstream tasks, showcasing strong recognition capabilities[14]. In the task of traffic sign recognition, leveraging CLIP’s powerful visual-language learning capability can enhance the understanding of sign semantic information, making it possible to achieve cross-region and environment recognition.
@@
In this work, to meet the requirements of cross-regional recognition, we first extract 46 mainstream and universal traffic sign categories from 10 existing datasets, creating a cross-regional traffic sign (CRTS) dataset. Based on the categories and feature distributions of traffic signs, we propose a prompt engineering scheme specifically designed for traffic signs. Regarding the TSCLIP model, we perform fine-tuning on the pre-trained CLIP model and introduce an adaptive dynamic weight ensembling (ADWE) fine-tuning scheme. Specifically, we dynamically integrate the results of each training iteration with the CLIP zero-shot model using adaptive factors. Ensuring the model maintains the generalization capabilities of the zero-shot model while learning new traffic sign knowledge. Our primary contributions are as follows:
@@
    We develop TSCLIP, the first comparative language-image model designed for traffic sign recognition, which achieves SOTA cross-region testing performance and outperforming mainstream benchmark models.
@@ II-A Traffic Sign Recognition
The advantage of deep learning methods lies in their ability to automatically learn and extract complex features from data, leading to higher accuracy [22]. [23] evaluated various CNNs and vision transformer models, showing that CNNs perform better in traffic sign classification. [24] proposed multi-scale CNN approaches that performed well in multiple datasets. [25] proposed a novel method using a limited common image set for traffic sign recognition, demonstrating excellent accuracy. [26] developed a lightweight CNN model that achieved near-perfect accuracy on the GTSRB dataset. [27] proposed a CNN model that achieved over 91% accuracy on an Indian dataset. Although deep learning methods can automatically learn features of different signs, their performance is often excellent only on the training scene. Once transferred to unseen scenarios, the robustness and generalization significantly decline.
@@ II-B3 Full Fine-Tuning (FFT)
FFT updates all parameters of the pre-trained model to adapt to specific tasks. The optimization objective involves updating the weights of both the image and text encoders. The loss function defined as:
@@ III-A1 Dataset Construction
In this study, we created a CRTS joint dataset based on mainstream open-source traffic sign datasets. To ensure regional diversity, we selected datasets from 10 different countries or regions[2, 3, 6, 9, 10, 11, 29, 28, 30, 31]. We extracted 46 commonly used categories by analyzing the distribution of similar categories and traffic sign regulations[32] across these datasets. Every dataset was cleaned, and all categories were standardized with corresponding names.
@@ III-A2 Difference of Cross-Regional Traffic Signs
Fig. 2 shows examples of four traffic sign categories (No Overtaking, No Parking, No Pedestrians, and Stop) in different regional contexts. Traffic sign patterns vary across countries and regions, influenced by local culture and traffic regulations. Some regions have unique sign patterns, such as the No Overtaking signs in China and America. Additionally, some signs incorporate local languages, as seen in the Stop signs from China, Iran, India, Turkey, and Russia. Therefore, cross-regional traffic sign recognition poses a significant challenge, requiring models to handle continuously changing patterns.
@@ III-B2 Refinement of Scenario Descriptions ($\mathbf{S_{i}}$)
The $\mathbf{S_{i}}$ encompasses four critical elements (a-d):
@@ III-B3 Traffic Sign Category and Rules ($\mathbf{T_{i}}$)
By following this structured approach, we create $n$ diverse and dynamic prompt templates $\{\mathbf{P_{1}},\mathbf{P_{2}},\ldots,\mathbf{P_{n}}\}$, ensuring a comprehensive representation of traffic signs in various scenarios. This method enhances the model’s ability to understand traffic signs by providing rich contextual information and explicit traffic rules, leading to more robust recognition.
@@ III-C TSCLIP Fine-Tuning Implementation
The core aim of the TSCLIP model’s robust fine-tuning framework (Fig. 3) is to combine the excellent generalization of the zero-shot model across all data distributions with the recognition capability of the fine-tuned model on traffic sign training data. This ensures the best of both worlds for specialized tasks and cross-regional generalization. Specifically, the fine-tuning framework consists of two main parts. Fig. 3(a) shows the initial fine-tuning of the TSCLIP using image data and corresponding prompts from the training samples, ensuring the fine-tuned model learns domain-specific knowledge of traffic signs. Fig. 3(b) illustrates the adaptive dynamic weight ensembling of the latest training weights with the CLIP zero-shot weight at the end of each training epoch, ensuring that the model retains a certain level of zero-shot generalization. Additionally, the robustness gains achieved by TSCLIP during the fine-tuning process do not incur extra computational costs during fine-tuning or inference.
@@ III-D1 Weight Ensembling
Interpolating model parameters is a classic idea in convex optimization[33]. Previous studies have shown that interpolation in the weight space can improve performance when models share part of the optimization trajectory [34]. Wise-FT is the first empirical study to explore the interpolation of non-convex models, specifically CLIP, from the perspective of distributional robustness. Fig. 3(c) illustrates Wise-FT’s ensembling method, which involves weight ensembling with the zero-shot model after the entire fine-tuning of the CLIP is completed. This weight ensembling method enhances robustness by forcibly injecting the weights of the zero-shot model into the final training results. Although experimental results show significant improvements, this approach overlooks the dynamic nature of CLIP weights during the fine-tuning process.
@@ III-D3 Adaptive Factor
In our dynamic weight ensembling method, since the weights obtained from each epoch are dynamically changing and the ensembled weights are further iteratively trained, a fixed mixing coefficient is not sufficient to handle the dynamic nature of the weights. The coefficient must adapt dynamically to the ever-changing training.
@@
This formula ensures that the ensemble proportion of the zero-shot model weights decreases smoothly over time while dynamically adjusting based on the relative performance of the trained model and the zero-shot model. This approach effectively balances the generalization and task-specific learning needs, leading to robust fine-tuning of the CLIP for traffic sign recognition across different regions.
@@ IV-A Experiment Setup
All models were trained in the PyTorch framework using the NVIDIA A100-PCIE-40GB GPUs. For training parameters, the CLIP-based models were set with a batch size of 512 and trained for 10 epochs with a learning rate of 0.00003. The classic classification model was trained based on the existing pre-trained model with a batch size of 128, trained for 100 epochs, and the learning rate was 0.0001.
@@ V Conclusion
To address the challenge of robust traffic sign recognition across regions and data distributions worldwide, we proposed the TSCLIP and CRTS benchmark dataset. Then, we developed a prompt engineering scheme that includes specific scene descriptions and corresponding rules, specifically made for traffic sign. The proposed ADWE method effectively combines fine-tuning model with zero-shot model, ensuring generalization to other environments while learning new traffic sign knowledge. In extensive cross-regional tests, TSCLIP significantly outperformed mainstream benchmark methods and achieved SOTA result compared to existing robust fine-tuning methods. The ablation experiments and visual analyses further validated and illustrated the effectiveness of our approach. Future research will involve collecting traffic sign data worldwide to build the foundation model, enabling general recognition of traffic signs across regions.